Solve the following for v.
h=4vt+16t^2

Answer by mananth(16946)

h=4vt+16t^2
add -16t^2
h-16t^2=4vt+16t^2-16t^2
h-16t^2=4vt
divide by 4t
h/4t - 16t^2/4t = v
h/4t -4t=v
...
m.ananth@hotmail.ca